OpenStudy (anonymous):

5s^2+19s-4 help please

OpenStudy (anonymous):

what is the question? factor?

OpenStudy (anonymous):

yes factor

OpenStudy (anonymous):

abc formula: -b/2a + or - sqrt (b^2-4ac)/2a

OpenStudy (anonymous):

(a+4)(5a-1)

OpenStudy (anonymous):

sorry in place of a u can write s

OpenStudy (anonymous):

find non-zero integers a b c d such that \[(a*s + b)(c*s + d) = 5s^2 + 19s - 4\]and \[a * c = 5\]\[b * d = -4\]\[b*c + a*d = 19\] 5 is prime so we've found \[a = 5\ \ \ and \ \ \ c = 1\] hence our problem is now to find b and d such that \[(5s + b)(s + d) = 5s^2 + 19s - 4\]where\[b*d = -4\]and \[b + 5d = 19\] it follows from \[b+5d = 19\]that \[b = 19 - 5d\]hence \[19d - 5d^2 = -4\]\[5d^2 - 19d -4 = 0\]solving gives us \[d = -\frac{1}{5} \ \ \ or \ \ \ d = 4\]we pick the integer solution hence \[b = -1\] therefore \[a = 5\]\[b = -1\]\[c = 1\]\[d = 4\] hence \[5s^2 +19s - 4 = (5s - 1)(s + 4)\] .... should have used that solution formula in the first step.